# Diffusion MRI with Machine Learning

### v3

@@ IV-D Microstructure mapping
There have been many efforts and much progress in developing bio-physical models that relate the measured dMRI signal to the microstructure of brain tissue [32, 31]. Advanced models rely on specialized measurement protocols and complex numerical optimization methods. The underlying estimation problem is typically nonlinear, sensitive to measurement noise and initialization, and may be unstable and computationally intensive [141, 142]. Accuracy and precision of model fitting is hard to verify and can significantly depend on the optimization algorithm used to compute the fitting [141]. To avoid local minima, some works have resorted to computationally intensive approaches such as grid search, multi-start methods, cascade optimization, and stochastic optimization techniques [143, 144, 142, 141]. There is a great interest in developing methods, such as those based on dictionary matching [145, 146], to reduce the computational time without compromising the estimation accuracy. A good example of such methods is AMICO [146], which reformulates the microstructure estimation equations as convex optimization problems that can be solved much faster.
@@ IV-D1 Potential advantages of machine learning methods
    With machine learning methods it is typically much easier to incorporate prior knowledge or additional information such as constraints on the parameters to be estimated or other MRI contrasts.
@@ IV-D2 Method design
Methodologically, these works have been dominated by neural networks, although other models such as random forests [45, 42, 60] and Bayesian estimation methods [153, 154] have also been used. The first work to propose deep neural networks for this purpose was q-space deep learning [16]. It reported accurate estimation of DKI parameters with 12 measurements and NODDI parameters with 8 measurements, resulting in a twelve-fold reduction in scan time. Many studies have followed [16] to develop neural network models for tissue microstructure mapping. This should not be surprising given that deep neural networks have emerged as the best-performing regression models in many applications [155]. Most non-neural-network methods either predate the recent surge of deep learning or do not include a rigorous comparison with neural networks.
Incorporation of population-estimated priors has been advocated by several studies [153, 156]. For estimation of NODDIDA (NODDI with Diffusivity Assessment), it was shown that estimating a multivariate Gaussian prior from 35 subjects significantly improved the prediction accuracy and robustness [153]. The authors concluded that incorporation of the prior reduced the ill-posedness of the estimation problem and made it possible to estimate this complex model from clinically-feasible measurements. Some works have suggested utilizing other inputs in addition to the dMRI measurements. For instance, one work used T1-weighted and T2-weighted images (registered to the dMRI data) to improve diffusion tensor estimation [46].
@@ IV-D3 Training procedures and data
Most studies have adopted an end-to-end training approach, where the input dMRI measurements are mapped to the target parameter of interest using a single (albeit deep) neural network. However, there are notable exceptions that are instructive. For DKI, Kerkela et al. [33] used the neural network predictions as input to a regularized non-linear least-squares method to compute the final values. This approach improved the estimation robustness of the standard method and reduced the probability of predicting implausible values. Other studies have also proposed to combine neural network estimators with conventional methods. For example for computing the NODDI parameters under very low signal to noise ratio (SNR), Gong et al. used neural networks to obtain good initial estimates, which were then used to obtain more accurate estimates using maximum likelihood estimation (MLE) [147]. The justification for this approach is that when SNR is low, MLE is prone to erroneous predictions because of its vulnerability to local minima. The predictions of the neural networks, on the other hand, although very fast, can suffer from small but significant biases. Hence, one can use the neural network to compute a good initial estimate, which can then be refined using an unbiased MLE estimator. A similar approach was used in [157], where a shallow neural net was used to estimate the isotropic volume fraction in the NODDI model, which was the used in a MLE formulation to compute the complete NODDI model from single-shell data.
For diffusion tensor estimation from six measurements, Liu et al. argue that directly mapping the measurements to the tensor is not optimal [151]. Given the measurements along six gradient directions, they employ a neural network to estimate the signal along additional directions. These measurements are subsequently used by a second network to estimate the diffusion tensor. Ye et al. [158, 159] have proposed methods based on sparse representations and deep neural networks. In one implementation, their method is a two-stage pipeline: the first stage uses an LSTM model to compute a sparse decomposition of the signal in a dictionary. An MLP computes the desired microstructure indices based on the sparse representation coefficients. This work has been in part inspired by the widely-used AMICO method [146], which decouples the estimation of NODDI and ActiveX [143] models into linear problems and solves them using sparsity-based optimization. Ye et al.’s experiments show that this method can estimate NODDI parameters with higher accuracy than standard estimation techniques. This method was further improved with the help of a separable spatial-angular dictionary for signal representation [160].
@@
The most common training approach has been supervised training, where the loss function to be minimized is computed based on the difference between the predicted and ground truth parameter values. However, there have been important and instructive exceptions. Kaandorp propose to train a neural network model in an unsupervised manner by using the predicted tissue parameters to predict the corresponding dMRI signal and optimizing the network weights to minimize the difference between the predicted and measured signal [52, 163]. They found that predictions of unsupervised methods had higher variability whereas predictions of supervised methods had a strong bias towards the mean of training data and were deceptively smooth [163]. Epstein et al. have analyzed the bias-variance trade-off in supervised and self-supervised microstructure estimation methods [164]. They argue that the reason for the lower bias of self-supervised methods is because they are based on the same optimization objective as MLE, while the high bias in supervised method is because they deviate from this objective and are based on a poor choice of training target. They show that using an MLE-computed parameter value as the estimation target can reduce the estimation bias of supervised methods.
@@ IV-D4 Evaluations and main results
It has been shown that models that leverage spatial correlations in a neighborhood (e.g., an image patch) around the voxel of interest can lead to more accurate estimation [161, 55, 165, 171]. Early works used CNNs to learn the spatial correlations. More recent studies haves relied on attention-based neural networks that are considered to be more effective in learning correlations [162, 22, 172]. Zheng et al. argue for the superiority of transformer architectures for tissue microstructure mapping in dMRI [162]. In order to enable effective implementation of these architectures with small datasets, they introduce additional computational modules. Specifically, they use a sparse representation stage to compute the signal embeddings in a dictionary using unfolded iterative hard thresholding.
@@
Although most studies have evaluated the new methods only in terms of estimation accuracy measures such as root mean square of the error (RMSE), several works have investigated the downstream use of the estimated tissue microstructure parameters. For example, Gibbons et al. showed that post‐stroke outcome prediction was the same for microstructure maps estimated with standard methods and those estimated with a neural network with 10-fold fewer measurements [161]. Aliotta et al. observed that deep learning-estimated FA maps resulted in more accurate delineation of the brain tumor boundaries in glioblastoma multiforme patients than with conventional methods [148]. Kaandorp et al. showed that intravoxel incoherent motion (IVIM) parameters computed by a deep learning method better predicted the chemoradiotherapy response of pancreatic cancer patients than a standard method [52]. Another work assessed the microstructure maps estimated by several deep learning methods in terms of their effectiveness in studying the impact of migraine on the brain white matter [173]. Specifically, FA and mean diffusivity (MD) values computed by different methods were compared using Tract-Based Spatial Statistics (TBSS) [174]. It was observed that, compared with a standard estimation technique, deep learning methods improved the true positive rate but also increased the false positive rate.
@@
For DTI estimation, several works have reported good reconstructions from the theoretical minimum of six measurements (Table II). DeepDTI [46] uses six diffusion-weighted measurements and a non-weighted (b0) measurement, T1-weighted, and T2-weighted images as input and computes the residuals with respect to high-quality measurements. The estimated high-quality measurements are then used to estimate the diffusion tensor with a standard estimation technique. Detailed analysis, including local assessment of DTI-derived parameters, tract-based analysis, and tractography show that DeepDTI achieves accurate estimation while reducing the number of measurements by a factor of at least 3-5. DeepDTI is based on a 3D FCN and requires that the six diffusion-weighted measurements be acquired along the optimal directions proposed in [175]. SuperDTI, proposed by Li et al. [55], also used an FCN for direct estimation of the diffusion tensor. It achieved accurate FA computation and tractography with six measurements and accurate MD computation with three measurements. Similar results were reported in [148, 171]. Using transformer networks, another study made a similar claim and further evaluated the method using tractography and quantitative structural connectivity assessment [51]. DiffNet [149] estimated FA and MD with 3-20 diffusion-weighted measurements and showed that the estimation accuracy was still higher than standard methods. DiffNet-estimated FA was more accurate for segmenting brain tumor volume, while also reducing the scan time.
@@ IV-E1 Method design
Unlike the scalar microstructural indices, FOD is a function of angle and, hence, it is rotation-sensitive. Elaldi et al. argue that the standard convolutional layers (which provide equivariance to planar translation) are insufficient for spherical signals in dMRI, where rotation equivariance is additionally needed [194]. Instead, they employ rotation-equivariant graph convolutions proposed in [210]. Furthermore, they opt for spherical harmonics of degree 20 (as opposed to degree 6 or 8 in standard methods [191]) to enable reconstruction of nearby FOD peaks. They show that, compared with the CSD baseline, their neural network achieves superior FOD estimation and better performance on downstream tasks such as tractography on phantom data and human brain scans. They have further extended this work by developing neural network layers that are equivariant to translations and grid reflections (as are standard FCNs) as well as to voxel and grid rotations [195]. They argue that these extra equivariance properties are needed to give the network the necessary spatio-spherical inductive biases to effectively learn to compute the FOD from data. Their results show more accurate FOD estimation, tractography, and brain tissue segmentation in dMRI compared with a range of state of the art methods. Related neural network architectures are reviewed in more detail in Section V-C.
@@ IV-E2 Training data
The majority of studies have used standard estimation methods such as CSD to generate target FODs for training and validation data [203, 204]. Typically, these works have aimed at matching the standard technique while using fewer measurements as input. It has been claimed that neural networks can use 6-20 measurements from a single shell to estimate the FOD estimated from multi-shell HARDI data computed with standard methods. Therefore, the validation has been based on comparison with standard methods. This is unlike the common validation approach for FOD estimation techniques, which has been primarily based on simulation [211]. However, with this approach, the trained model will inevitably inherit some of the shortcomings of the standard method that is used to generate the target FOD [194, 183, 184]. Moreover, the assessments can only tell us how close the new machine learning method is to the standard technique.
@@ IV-G Delineation of white matter tracts
BrainSegNet [270] Bidirectional LSTM Tractogram 8 The method works in two steps, in the first step classifying the white matter versus the gray matter and in the second step segmenting the tracts. The method achieves streamline classification accuracy of $>96\%$ and recall of $>73\%$, but it is tested on three subjects only.
@@ V-A1 Validation on abnormal/pathological data
For tissue microstructure mapping, some studies have validated the new methods on selected pathologies such as MS [16] and stroke [55]. One study trained a model on healthy brain data to jointly estimate the DTI, DKI, and multi-compartment models and tested the model on stroke patients [165]. Results showed that the model accuracy on pathological tissue suffered more for multi-compartment models (e.g., NODDI) than for DTI. Similarly, a few studies have validated tract-specific analysis methods on clinical data. For example, one study [366] showed that automatic tract segmentation with TractSeg [17] was more accurate in detecting microstructural changes due to traumatic brain injury than a method based on manual tract segmentation. Li at al. [310] validated their joint tract segmentation and registration method on a clinical population. Although most of the published results on abnormal brains are positive and encouraging, they are not sufficient. Validation of deep learning methods on abnormal brain data is especially important because these methods are highly complex and difficult to interpret, and because the generalizability of these methods to unseen data is hard to predict [367]. As an example, for tissue microstructure mapping one study has found that training with a loss function in terms of error in microstructure estimation may lead to better generalizability to pathological brains than loss functions in terms of signal reconstruction error [178]. Other factors may also influence the generalizability of machine learning methods to pathological test data, but there has been little work to investigate this topic.
@@ V-A2 Validation on external data
A number of works have reported successful application of trained machine learning models on data from other scanners/centers without any data harmonization or re-training [162]. Other studies have suggested that when dMRI intensities or subject demographics deviate from training data, intensity normalization, data harmonization, or model re-training/fine-tuning may be necessary [16, 368]. However, data harmonization in dMRI is a complex problem of its own as mentioned on in Section IV-C. Moreover, there may be a tendency in the published works not to report discouraging results when it comes to validation on external data. Given the significant diversity across datasets and applications and strong dependence of machine learning methods on training data distribution, model settings, and design choices, more extensive validations are needed to assess the true advantages of machine learning methods in real-world settings with heterogeneous data.

### v5

@@ IV-D Microstructure mapping
There have been many efforts and much progress in developing bio-physical models that relate the measured dMRI signal to the microstructure of brain tissue [32, 31]. Advanced models rely on specialized measurement protocols and complex numerical optimization methods. The underlying estimation problem is typically nonlinear, sensitive to measurement noise and initialization, and may be unstable and computationally intensive [141, 142]. Accuracy and precision of model fitting is hard to verify and can significantly depend on the optimization algorithm used to fit the model to the measurements [141]. To avoid local minima, some works have resorted to computationally intensive approaches such as grid search, multi-start methods, cascade optimization, and stochastic optimization techniques [143, 144, 142, 141]. There is a great interest in developing methods, such as those based on dictionary matching [145, 146], to reduce the computational time without compromising the estimation accuracy. An example of such methods is AMICO [146], which reformulates the microstructure estimation equations as convex optimization problems that can be solved much faster.
@@ IV-D1 Potential advantages of machine learning methods
    With machine learning methods it is typically much easier to incorporate prior knowledge or additional information. For example, it is generally easy to impose constraints on the parameter values to be estimated and to include other MRI contrasts as input to machine learning models.
@@ IV-D2 Method design
Methodologically, these works have been dominated by neural networks, although other models such as random forests [45, 42, 60] and Bayesian estimation methods [153, 154] have also been used. The first work to propose deep neural networks for this purpose was the “q-space deep learning” method [16]. It reported accurate estimation of DKI parameters with 12 measurements and NODDI parameters with 8 measurements, resulting in a twelve-fold reduction in scan time. Many studies have followed [16] to develop neural network models for tissue microstructure mapping. This should not be surprising given that deep neural networks have emerged as the best-performing regression models in many applications [155]. Most non-neural-network methods either predate the recent surge of deep learning or do not include a rigorous comparison with neural networks.
Incorporation of population-estimated priors has been advocated by several studies [153, 156]. For estimation of NODDIDA (NODDI with Diffusivity Assessment), it was shown that estimating a multivariate Gaussian prior from 35 subjects significantly improved the prediction accuracy and robustness [153]. The authors concluded that incorporation of the prior reduced the ill-posedness of the estimation problem and made it possible to estimate this complex model from clinically-feasible measurements. Some works have suggested utilizing other inputs in addition to the dMRI measurements. For instance, one work has used T1-weighted and T2-weighted images (registered to the dMRI data) to improve diffusion tensor estimation [46].
@@ IV-D3 Training procedures and data
Most studies have adopted an end-to-end training approach, where the input dMRI measurements are mapped to the target parameter of interest using a single (albeit deep) neural network. However, there are notable exceptions that are instructive. For DKI, Kerkela et al. [33] used the neural network predictions as input to a regularized non-linear least-squares method to compute the final values. This approach improved the estimation robustness of the standard method and reduced the probability of predicting implausible values. Other studies have also proposed to combine neural network estimators with conventional methods. For example for computing the NODDI parameters under very low signal to noise ratio (SNR), Gong et al. used neural networks to obtain good initial estimates, which were then used to obtain more accurate estimates using maximum likelihood estimation (MLE) [147]. The justification for this approach is that when the SNR is low, MLE is prone to erroneous predictions because of its vulnerability to local minima. The predictions of the neural networks, on the other hand, although very fast, can suffer from relatively small but significant biases. Hence, one can use the neural network to compute a good initial estimate, which can then be refined using an unbiased MLE estimator. A similar approach was proposed in [157], where a shallow neural network was applied to estimate the isotropic volume fraction in the NODDI model, which was subsequently used in an MLE formulation to compute the complete NODDI model from single-shell data.
For diffusion tensor estimation from six measurements, Liu et al. argue that directly mapping the measurements to the diffusion tensor is not optimal [151]. Given the measurements along six gradient directions, they employ a neural network to estimate the signal along additional directions. These measurements are subsequently used by a second network to estimate the diffusion tensor. Ye et al. [158, 159] have proposed methods based on sparse representations and deep neural networks. In one implementation, their method is a two-stage pipeline: the first stage uses an LSTM model to compute a sparse decomposition of the signal in a dictionary. An MLP computes the desired microstructure indices based on the sparse representation coefficients. This work has been in part inspired by the widely-used AMICO method [146], which decouples the estimation of NODDI and ActiveX [143] models into linear problems and solves them using sparsity-based optimization. Ye et al.’s experiments show that this method can estimate NODDI parameters with higher accuracy than standard estimation techniques. This method was further improved with the help of a separable spatial-angular dictionary for signal representation [160].
@@
The most common training approach has been supervised training, where the loss function to be minimized is computed based on the difference between the predicted and ground truth parameter values. However, there have been important and instructive exceptions. Kaandorp propose to train a neural network model in an unsupervised manner by using the predicted tissue parameters to predict the corresponding dMRI signal and optimizing the network weights to minimize the difference between the predicted and measured signal [52, 163]. They found that predictions of the unsupervised methods had higher variability compared with the supervised methods. Predictions of the supervised methods, on the other hand, displayed a strong bias towards the mean of the training data and were deceptively smooth [163]. Epstein et al. have analyzed the bias-variance trade-off in supervised and self-supervised microstructure estimation methods [164]. They argue that the reason for the lower bias of self-supervised methods is because they are based on the same optimization objective as MLE, while the higher prediction bias of supervised methods is because they deviate from this objective and are based on a poor choice of training target. They show that using an MLE-computed parameter value as the estimation target can reduce the estimation bias of supervised methods.
@@ IV-D4 Evaluations and main results
Furthermore, it has been shown that models that leverage spatial correlations in a neighborhood (e.g., an image patch) around the voxel of interest can lead to more accurate estimation [161, 55, 165, 171]. Early works used CNNs to learn the spatial correlations. More recent studies have relied on attention-based neural networks that are considered to be more effective in learning correlations [162, 22, 172]. Zheng et al. argue for superiority of transformer architectures for tissue microstructure mapping in dMRI [162]. In order to enable effective implementation of these architectures with small datasets, they introduce additional computational modules. Specifically, they use a sparse representation stage to compute the signal embeddings in a dictionary using unfolded iterative hard thresholding.
@@
Although most studies have evaluated the new methods only in terms of estimation accuracy metrics such as root mean square of the error (RMSE), several works have investigated downstream use of the estimated tissue microstructure parameters. For example, Gibbons et al. showed that post‐stroke outcome prediction was the same for microstructure maps estimated with standard methods and those estimated with a neural network with 10-fold fewer measurements [161]. Aliotta et al. observed that deep learning-estimated FA maps resulted in more accurate delineation of the brain tumor boundaries in glioblastoma multiforme patients than with conventional methods [148]. Kaandorp et al. showed that intravoxel incoherent motion (IVIM) parameters computed by a deep learning method better predicted the chemoradiotherapy response of pancreatic cancer patients than a standard method [52]. Another work assessed the microstructure maps estimated by several deep learning methods in terms of their effectiveness in studying the impact of migraine on the brain white matter [173]. Specifically, FA and mean diffusivity (MD) values computed by different methods were compared using Tract-Based Spatial Statistics (TBSS) [174]. It was observed that, compared with a standard estimation technique, deep learning methods improved the true positive rate but also increased the false positive rate.
@@
For DTI estimation, several works have reported good reconstructions from the theoretical minimum of six measurements (Table II). DeepDTI [46] uses six diffusion-weighted measurements and a non-weighted (b0) measurement, T1-weighted, and T2-weighted images as input and computes the residuals with respect to high-quality measurements. The estimated high-quality measurements are then used to estimate the diffusion tensor with a standard estimation technique. Detailed analysis, including local assessment of DTI-derived parameters, tract-based analysis, and tractography show that DeepDTI achieves accurate estimation while reducing the number of measurements by a factor of at least three. DeepDTI is based on a 3D FCN and requires that the six diffusion-weighted measurements be acquired along the optimal directions proposed in [175]. SuperDTI, proposed by Li et al. [55], also used an FCN for direct estimation of the diffusion tensor. It achieved accurate FA computation and tractography with six measurements and accurate MD computation with three measurements. Similar results were reported in [148, 171]. Using transformer networks, another study made a similar claim and further evaluated the method using tractography and quantitative structural connectivity assessment [51]. DiffNet [149] estimated FA and MD with 3-20 diffusion-weighted measurements and showed that the estimation accuracy was higher than standard methods. DiffNet-estimated FA was more accurate for segmenting brain tumor volume, while also reducing the scan time.
@@ IV-E1 Method design
Unlike the scalar microstructural indices, FOD is a function of angle and, hence, it is rotation-sensitive. Elaldi et al. argue that the standard convolutional layers (which provide equivariance to planar translation) are insufficient for spherical signals in dMRI, where rotation equivariance is additionally needed [194]. Instead, they employ rotation-equivariant graph convolutions proposed in [210]. Furthermore, they opt for spherical harmonics of degree 20 (as opposed to degree 6 or 8 in standard methods [191]) to enable reconstruction of nearby FOD peaks. They show that, compared with the CSD baseline, their neural network achieves superior FOD estimation and better performance on downstream tasks such as tractography on phantom data and human brain scans. They have further extended this work by developing neural network layers that are equivariant to translations and grid reflections (as are standard FCNs) as well as to voxel and grid rotations [195]. They argue that these extra equivariance properties are needed to give the network the necessary spatio-spherical inductive biases to effectively learn to compute the FOD from data. Their results show more accurate FOD estimation, tractography, and brain tissue segmentation in dMRI compared with a range of state of the art methods. Related neural network architectures are reviewed in more detail in Section V-C of this paper.
@@ IV-E2 Training data
The majority of studies have used standard estimation methods such as CSD to generate target FODs for training and validation data [203, 204]. Typically, these works have aimed at matching the standard technique while using fewer measurements as input. It has been claimed that neural networks can compute the FOD using only 6-20 measurements from a single shell to match the FOD computed with standard methods using multi-shell HARDI data. Therefore, the validation has been based on comparison with standard methods. This is unlike the common validation approach for FOD estimation techniques, which has been primarily based on simulation [211]. However, with this approach, the trained model will inevitably inherit some of the shortcomings of the standard method that is used to generate the target FOD [194, 183, 184]. Moreover, the assessments can only tell us how close the new machine learning method is to the standard technique.
@@ IV-G Delineation of white matter tracts
BrainSegNet [270] Bidirectional LSTM Tractogram 8 The method works in two steps, in the first step classifying the white matter versus the gray matter and in the second step segmenting the tracts. The method achieves streamline classification accuracy of $96\%$ and recall of $73\%$, but it is tested on three subjects only.
@@ IV-G3 Superficial white matter analysis methods
Li et al. developed a deep learning–based framework for precise alignment of diffusion tensor images of the fetal brain [341]. This is a particularly challenging problem due to the low quality and scarcity of fetal imaging data, the rapid structural changes that occur during gestation, and the limited availability of anatomical landmarks to guide registration. The proposed method, FetDTIAlign, performs both affine and deformable registration. The affine registration module adopts an iterative strategy: the output of each alignment step is recursively fed back into the model to progressively refine accuracy. For deformable registration, FetDTIAlign leverages segmentation masks of 60 white matter tracts, which serve as anatomical constraints to improve robustness and precision. Extensive evaluations demonstrated that FetDTIAlign outperformed both conventional and deep learning–based registration approaches across a wide range of gestational ages (23–36 weeks). Such tools are urgently needed, as diffusion MRI is increasingly applied to the study of fetal and neonatal brain development, yet most existing registration methods have been designed for the adult brain and fail to address the unique challenges of early brain imaging.
@@ V-A1 Validation on abnormal/pathological data
For tissue microstructure mapping, some studies have validated the new methods on selected pathologies such as MS [16] and stroke [55]. One study trained a model on healthy brain data to jointly estimate the DTI, DKI, and multi-compartment models and tested the model on stroke patients [165]. Results showed that the model accuracy on pathological tissue suffered more for multi-compartment models (e.g., NODDI) than for DTI. Similarly, a few studies have validated tract-specific analysis methods on clinical data. For example, one study [367] showed that automatic tract segmentation with TractSeg [17] was more accurate in detecting microstructural changes due to traumatic brain injury than a method based on manual tract segmentation. Li at al. [310] validated their joint tract segmentation and registration method on a clinical population.
Although most published results on abnormal brains are positive and encouraging, they are not sufficient. Validation of deep learning methods on abnormal brain data is especially important because these methods are highly complex and difficult to interpret, and because the generalizability of these methods to unseen data is hard to predict [368]. As an example, for tissue microstructure mapping one study has found that training with a loss function in terms of error in microstructure estimation may lead to better generalizability to pathological brains than loss functions in terms of signal reconstruction error [178]. Other factors may also influence the generalizability of machine learning methods to pathological test data, but there has been little work to investigate this topic.
@@ V-A2 Validation on external data
A number of works have reported successful application of trained machine learning models on data from other scanners/centers without any data harmonization or re-training [162]. Other studies have suggested that when dMRI intensities or subject demographics deviate from training data, intensity normalization, data harmonization, or model re-training/fine-tuning may be necessary [16, 369]. However, data harmonization in dMRI is a complex problem of its own as mentioned above in Section IV-C. Moreover, there may be a tendency in the published works not to report discouraging results when it comes to validation on external data. Given the significant diversity across datasets and applications and strong dependence of machine learning methods on training data distribution, model settings, and design choices, more extensive validations are needed to assess the true advantages of machine learning methods in real-world settings with heterogeneous data.